# Confidence Improves Self-Consistency in LLMs

### v1

@@ 2 Notations
[…]
@@ 4.1 Confidence Extraction Methods
[…]
@@
[…]
@@ 4.2 Datasets
  • MMLU-Pro Wang et al. (2024c): A more challenging version of the Multitask Language Understanding (MMLU) benchmark, testing language models’ general knowledge and reasoning abilities with a wide range of topics such as science and history. We randomly sampled 5K questions.
@@ 4.4 Metrics
    100×(1CISC budget# Comparable SC $100\times\left(1-\frac{\text{CISC budget}}{\text{\# Comparable SC responses}}\right)$
@@
    100×(CISC AccSC $100\times\left(\frac{\text{CISC Acc}}{\text{SC Acc}}-1\right)$
@@ 4.5 Temperature Scaling
As discussed in §3, CISC re-scales the confidence values using a softmax transformation, parameterized by a temperature $T>0$. We tune the temperature separately for each model and confidence extraction method using a 10% held-out set, aggregated across all four datasets (§4.2). More details and the optimal temperature values for each configuration are in appendix D.
@@ 5 Main Results
Confidence Method Cost Reduction @ 10
@@ 6 Within-Question Confidence Evaluation
Confidence Method ECE-t $\downarrow$ Brier-t $\downarrow$ WQD $\uparrow$ CISC Cost Reduction $\uparrow$
@@ 8 Related Work
There is a long line of work on deriving confidence measures from LLMs. Popular approaches use the agreement across multiple samples Kuhn et al. (2023); Manakul et al. (2023); Tian et al. (2023a); Lyu et al. (2024), the model’s internal representations Azaria and Mitchell (2023); Burns et al. (2022) or directly prompting the model to verbalize its confidence Tian et al. (2023b); Kadavath et al. (2022). All papers in this line of work focused on fact-seeking tasks, so confidence is typically derived based on the final answer alone. To the best of our knowledge, our work is the first to apply these approaches to scoring the entire reasoning path.
@@
Numerous attempts Chen et al. (2024) have been made to reduce SC computational overhead while maintaining quality. However, none have matched the widespread adoption of self-consistency. This can be largely attributed to several limitations: (1) a trade-off where throughput is reduced while latency increases, for example by sampling chains sequentially until reaching a certain condition Li et al. (2024b) or running expensive LLM calls instead of the cheap majority voting Yoran et al. (2023), (2) the need for manual feature crafting and tuning tailored to each dataset Wan et al. (2024), (3) promising results on specialized setups Wang et al. (2024a) which did not generalize to standard benchmarks (Table 9), and (4) as highlighted by Huang et al. (2023), many of the more sophisticated methods that appear promising actually don’t outperform self-consistency when evaluated with a thorough analysis of inference costs. Our approach is different in that CISC adds minimal complexity to self-consistency, and improves throughput without compromising latency.
@@ 9 Discussion
We encourage future research to explore the integration of model self-confidence into more sophisticated reasoning frameworks like Tree of Thoughts Yao et al. (2024) or Graph of Thoughts Besta et al. (2024), believing that harnessing this inherent capability can further boost performance. Another promising avenue is training models to produce more accurate intrinsic or verbal confidence Lin et al. (2022); Chaudhry et al. (2024), which would directly improve CISC and related methods. Conversely, CISC and WQD can be used to assess the impact of advancements in confidence generation.
@@ 10 Limitations
The preferred CISC approach calculates P(True) (as described in §4.1) by examining the model’s assigned probability to the verbal confidence token. This method is available in both popular open-weight frameworks (e.g., Hugging-Face (2024a)) and closed-weight frameworks (e.g., OpenAI (2025)). However, this feature may not be universally available across all frameworks.
@@ Appendix A Quantitative example from §3
Figure 5: The relationship between the number of samples (x-axis) and the accuracy of majority vote over these samples (y-axis), for two different hypothetical cases: (blue) Each sample receives an equal weight in majority voting, and (orange) Correct answers are assigned double the weight of incorrect ones. Adding this additional weighting information translated into $4X$ reduction in the number of samples required for the majority vote to reach 90% accuracy.
@@ Appendix B Prompting Techniques
MMLU-Pro You will be given a single-choice question. Answer the question by selecting the letter of the best fitting option. [General Instructions] The answer MUST ALWAYS be the letter of one of the available options; it CANNOT be "None of the Above".
@@
BBH (no options) You will be given a question and your goal is to answer it correctly. [General Instructions]
BBH (with options) You will be given a question and your goal is to answer it correctly. [General Instructions] Select the letter of the best fitting option. The answer CANNOT be "None of the Above".
GSM8K You will be given a question and your goal is to answer it correctly. [General Instructions]
@@
Verbal 0-100 Now I will rate my confidence in the proposed answer on a scale of 0-100. Proposed confidence: (
Verbal Binary Now I will rate my confidence in the proposed answer as either 0 or 1. Proposed confidence: (
@@ Appendix C Additional Results
We also add Figure 6 featuring additional graphs similar to Figure 1, but with all the confidence methods.
@@ Appendix D Temperature Scaling Results
As discussed in §4.5, a single optimal temperature, $T^{\star}$, was determined for each model and confidence extraction method by using a 10% held-out set, aggregated across all datasets. Fitting is done using grid search on values between $10^{-4}$ and $10^{4}$. The temperatures for each configuration are presented in Figure 7. As can be seen, each of the confidence extraction method work with a different temperature magnitude because it produce confidence values on a different scale.
@@ Appendix E Qualitative Appendix
Human evaluators were asked to identify logical patterns in the LLMs’ reasoning paths that reduced the evaluators’ confidence in the correctness of the LLMs’ answers. Importantly, the MMLU dataset requires significant domain knowledge and unspecialized humans achieved only 34.5% accuracy Hendrycks et al. (2020), compared to a random baseline of 25%. The MMLU-pro dataset is based on the MMLU dataset, but is considered much harder. This means that our evaluators, which lacked specialized knowledge, could not easily how to solve each question. Instead, we instructed them to focus on identifying low-quality reasoning errors in the responses of the LLMs. This approach aligns with findings from a prior analysis on GPT-4o Wang et al. (2024b), which attributed 39% of its errors to reasoning flaws that do not rely on specialized domain knowledge.
@@
No choice The model arrives at a solution which is not present in the list of available options. This can include case where a mathematical answer significantly diverging from all options, answers that are only partially correct, or the elimination of all options as part of the reasoning process. 38% 13% "… After reviewing the options, it’s clear that none of them perfectly fit the requirements. However, the closest correct option is (A), which only has a minor error in calculating the remaining inches. Proposed answer: (A)"
Incomplete Calculations The model begins to solve the problem but does not complete the full calculation, often due to the lack of necessary data. For example, when attempting to compute acceleration, the absence of mass data prevents an exact and full calculation. 22% 2% "…**Calculate Heat Flow:** q" = h * (T-surface - T-air) **Note:** Without the actual values for air density, viscosity, and thermal conductivity at 68°F, we cannot perform the precise calculations. Proposed answer: (C)."
Multiple candidates The model explores several plausible solutions without identifying a definitive "correct" one. This occurs when the model solves a problem generally, relying on estimations rather than concrete data, resulting in a range of potential answers. 11% 16% "… 2.**Identify Buddhist Thinkers:** The options list several prominent Buddhist figures from various traditions… 4. **Most Prominent:** The Dalai Lama and Thich Nhat Hanh stand out for their consistent emphasis on self-sacrifice in their teachings and actions. Proposed answer: (I)"
@@
General Instructions Evaluate the LLMs’ reasoning paths, looking for logical inconsistencies or errors that lower your confidence in their conclusions. Because the questions are very difficult, even for experts, your task is to identify general reasoning flaws, not to assess the correctness of the final answers themselves. Examples: Incorrect Assumption: The model assumes something without justification Missing Step: The model skips a crucial step in the reasoning process Contradiction: The model states both A and not-A
@@
LLM Output [Pre-filled - The LLM output for the given question]

### v2

@@ 1 Introduction
  • We open-source our implementation of CISC to facilitate further research [^1] Our code is available at https://github.com/google-research/google-research/tree/master/cisc..
@@ 2 Notations
We consider an auto-regressive language model $M$ with parameters $\theta$. We use $p_{\theta}(\cdot|x)$ to denote $M$’s distribution over the next token given the provided context $x$. Given a question $q$ (e.g., “Jane had 4 apples and ate half of her apples. How many apples she has now?”), we denote the model’s response as $(\textbf{r},\textbf{a})$, where a is the answer (e.g., “2”) and r is a reasoning path (or chain-of-thought), a sequence of logical steps supposedly leading up to this answer (e.g., “If Jane ate half her apples, this means she ate 2 apples. 4 minus 2 is 2.”).
@@ 4.1 Confidence Extraction Methods
  • Response Probability Wang et al. (2022): The confidence in a response $(\textbf{r},\textbf{a})$ is taken to be the model’s (length-normalized) probability of generating $(\textbf{r},\textbf{a})=(x_{1},\dots,x_{n})$ given the question:
@@
  • Verbal Confidence Lin et al. (2022): After sampling $(\textbf{r},\textbf{a})$ from the model, we prompt it to rate its confidence in its previously generated output. We implement two variants: (1) Verbal Binary instructs the model to output either 0 or 1, and (2) Verbal 0-100 instructs the model to output a score on a scale of 0-100.
  • P(True) Kadavath et al. (2022): We prompt the model to rate its confidence in $(\textbf{r},\textbf{a})$ in binary format (either 0 or 1), and compute the probability that the model assigns to the token $1$.
@@ 4.2 Datasets
  • MMLU-Pro Wang et al. (2024d): A more challenging version of the Multitask Language Understanding (MMLU) benchmark, testing language models’ general knowledge and reasoning abilities with a wide range of topics such as science and history. We randomly sampled 5K questions.
@@ 4.4 Metrics
    $100\times\left(1-\frac{\text{CISC budget}}{\text{\# Comparable SC responses}}\right)$
@@
    $100\times\left(\frac{\text{CISC Acc}}{\text{SC Acc}}-1\right)$
@@ 4.5 Temperature Scaling
As discussed in §3, CISC re-scales the confidence values using a softmax transformation, parameterized by a temperature $T>0$. We tune the temperature separately for each model and confidence extraction method using a 10% held-out set, aggregated across all four datasets (§4.2). The fact that CISC only employees a single dataset-agnostic hyper-parameter, makes the tuning process light-weight and robust. More details and the optimal temperature values for each configuration are in appendix D.
@@ 5 Main Results
Confidence Method
Cost Reduction @ 10
@@ 6 Within-Question Confidence Evaluation
Confidence
Method
ECE-t $\downarrow$ Brier-t $\downarrow$
WQD $\uparrow$
CISC Cost Reduction $\uparrow$
@@ 8 Related Work
There is a long line of work on deriving confidence measures from LLMs. Popular approaches use the agreement across multiple samples Kuhn et al. (2023); Manakul et al. (2023); Tian et al. (2023a); Lyu et al. (2024); Gekhman et al. (2024), the model’s internal representations Azaria and Mitchell (2023); Burns et al. (2022); Orgad et al. (2025) or directly prompting the model to verbalize its confidence Tian et al. (2023b); Kadavath et al. (2022). All papers in this line of work focused on fact-seeking tasks, so confidence is typically derived based on the final answer alone. To the best of our knowledge, our work is the first to apply these approaches to scoring the entire reasoning path.
@@
Numerous attempts Chen et al. (2024) have been made to reduce SC computational overhead while maintaining quality. However, none have matched the widespread adoption of self-consistency. This can be largely attributed to several limitations: (1) a trade-off where throughput is reduced while latency increases, for example by sampling chains sequentially (instead of in parallel) until reaching a certain condition Aggarwal et al. (2023); Li et al. (2024b); Wang et al. (2024b) or running expensive LLM calls instead of the cheap majority voting Yoran et al. (2023), (2) the need for manual feature crafting and tuning tailored to each dataset Wan et al. (2024), (3) promising results on specialized setups Wang et al. (2024a) which did not generalize to standard benchmarks (Table 9), and (4) as highlighted by Huang et al. (2023), many of the more sophisticated methods that appear promising actually don’t outperform self-consistency when evaluated with a thorough analysis of inference costs. Our approach is different in that CISC adds minimal complexity to self-consistency, and still allows parallel sampling which enables to improve throughput without compromising latency, a crucial requirement for many applications.
@@ 9 Discussion
We encourage future research to explore the integration of model self-confidence into more sophisticated reasoning frameworks like Tree of Thoughts Yao et al. (2024) or Graph of Thoughts Besta et al. (2024), believing that harnessing this inherent capability can further boost performance. Another promising avenue is training models to produce more accurate intrinsic or verbal confidence Lin et al. (2022); Chaudhry et al. (2024), which would directly improve CISC and related methods. For instance, recent evidence suggests that a better signal can be derived from the model’s internal states, even outperforming P(True) Gekhman et al. (2025). Conversely, CISC and WQD can be used to assess the impact of advancements in confidence generation.
@@ 10 Limitations
The preferred CISC approach calculates P(True) (as described in §4.1) by examining the model’s assigned probability to the verbal confidence token. This method is available in both popular open-weights frameworks (e.g., Hugging-Face (2024a)) and closed-weights frameworks (e.g., OpenAI (2025)). However, this feature may not be universally available across all frameworks.
@@ Appendix A Quantitative example from §3
Figure 5: The relationship between the number of samples (x-axis) and the accuracy of majority vote over these samples (y-axis), for two different hypothetical cases sampled from a Binomial distribution: (blue) Each sample receives an equal weight in majority voting, and (orange) Correct answers are assigned double the weight of incorrect ones. Adding this additional weighting information translated into $4X$ reduction in the number of samples required for the majority vote to reach 90% accuracy.
@@ Appendix B Prompting Techniques
MMLU-Pro
You will be given a single-choice question. Answer the question by selecting the letter of the best fitting option.
[General Instructions]
The answer MUST ALWAYS be the letter of one of the available options; it CANNOT be "None of the Above".
MATH
@@
[General Instructions]
BBH
(no options)
You will be given a question and your goal is to answer it correctly.
[General Instructions]
BBH
(with options)
You will be given a question and your goal is to answer it correctly.
[General Instructions]
Select the letter of the best fitting option. The answer CANNOT be "None of the Above".
GSM8K
You will be given a question and your goal is to answer it correctly.
[General Instructions]
@@
Verbal 0-100
Now I will rate my confidence in the proposed answer on a scale of 0-100. Proposed confidence: (
Verbal Binary
Now I will rate my confidence in the proposed answer as either 0 or 1. Proposed confidence: (
@@ Appendix C Additional Results
We also add Figures 6, 7 featuring additional graphs similar to Figure 1, but with all the confidence methods.
@@
Figure 7: Comparison between different confidence extraction methods using Mistral 123B model and four datasets (§4.2). CISC with P(True) outperforms Self-Consistency in all 4 graphs and is the best of all the CISC variants in 3 graphs.
@@ Appendix D Temperature Scaling Results
As discussed in §4.5, a single optimal temperature, $T^{\star}$, was determined for each model and confidence extraction method by using a 10% held-out set, aggregated across all datasets. Fitting is done using grid search on 80 evenly spaced values ranging from $10^{-4}$ to $10^{4}$. This was a light-weight process, only taking a few minutes on a standard desktop since no LLM re-runs were necessary. The temperatures for each configuration are presented in Figure 8. As can be seen, each of the confidence extraction method work with a different temperature magnitude because it produce confidence values on a different scale.
@@ Appendix E Qualitative Appendix
Human evaluators were asked to identify logical patterns in the LLMs’ reasoning paths that reduced the evaluators’ confidence in the correctness of the LLMs’ answers. Importantly, the MMLU dataset requires significant domain knowledge and unspecialized humans achieved only 34.5% accuracy Hendrycks et al. (2020), compared to a random baseline of 25%. The MMLU-pro dataset is based on the MMLU dataset, but is considered much harder. This means that our evaluators, which lacked specialized knowledge, could not easily how to solve each question. Instead, we instructed them to focus on identifying low-quality reasoning errors in the responses of the LLMs. This approach aligns with findings from a prior analysis on GPT-4o Wang et al. (2024c), which attributed 39% of its errors to reasoning flaws that do not rely on specialized domain knowledge.
@@
No choice
The model arrives at a solution which is not present in the list of available options. This can include case where a mathematical answer significantly diverging from all options, answers that are only partially correct, or the elimination of all options as part of the reasoning process.
38% 13%
"… After reviewing the options, it’s clear that none of them perfectly fit the requirements. However, the closest correct option is (A), which only has a minor error in calculating the remaining inches. Proposed answer: (A)"
Incomplete
Calculations
The model begins to solve the problem but does not complete the full calculation, often due to the lack of necessary data. For example, when attempting to compute acceleration, the absence of mass data prevents an exact and full calculation.
22% 2%
"…**Calculate Heat Flow:** q" = h * (T-surface - T-air) **Note:** Without the actual values for air density, viscosity, and thermal conductivity at 68°F, we cannot perform the precise calculations. Proposed answer: (C)."
Multiple
candidates
The model explores several plausible solutions without identifying a definitive "correct" one. This occurs when the model solves a problem generally, relying on estimations rather than concrete data, resulting in a range of potential answers.
11% 16%
"… 2.**Identify Buddhist Thinkers:** The options list several prominent Buddhist figures from various traditions… 4. **Most Prominent:** The Dalai Lama and Thich Nhat Hanh stand out for their consistent emphasis on self-sacrifice in their teachings and actions. Proposed answer: (I)"
@@
General
Instructions
Evaluate the LLMs’ reasoning paths, looking for logical inconsistencies or errors that lower your confidence in their conclusions. Because the questions are very difficult, even for experts, your task is to identify general reasoning flaws, not to assess the correctness of the final answers themselves. Examples: Incorrect Assumption: The model assumes something without justification Missing Step: The model skips a crucial step in the reasoning process Contradiction: The model states both A and not-A
@@
LLM
Output
[Pre-filled - The LLM output for the given question]
@@ Appendix F Compute
We consider 30 samples to be a substantial sample size. In more practical scenarios, we anticipate practitioners would likely use a smaller number of samples. As illustrated in Figures 1 and 6, the improvement curves show a logarithmic shape. On average, across all models and datasets, we found that just 13 responses were sufficient to achieve 90% of the maximum effect observed with 30 responses.